OpenStudy (venesa):

whether cos2acos2b equal to 1?

OpenStudy (anonymous):

\[(\cos ^{2}a -\sin ^{2}a)(\cos ^{2}b-\sin ^{2}b)\]....

OpenStudy (venesa):

whether it is equal to 1 ??

OpenStudy (anonymous):

simplfying from the step i gave you it will result to zero

OpenStudy (venesa):

ok